# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v2

@@ 3.4. Crowdsourcing Robot Behaviors
For each of our developed scenarios, we hoped to capture participant preferences for the robot’s behavior in response to its collection of user data. We did not, however, have a set of standardized robot response options that we could present to participants and, later, our selected LLMs. Therefore, we conducted a second online study to crowdsource these options to robot responses.
@@ 3.4.1. Procedure
In this study, each participant was assigned one of two conditions. In the first condition, participants were told to imagine themselves going about their daily routines at home when their robot records private information about them. In the second condition, participants were told to imagine the same scenario, however, they are provided additional hypothetical detail to ground their responses (i.e., “This scenario may involve you alone, your interactions with other people, or interactions with the robot, and may occur in shared or private spaces within the home.”). We created these two conditions to collect responses that were either entirely unbiased or grounded in context. Participants were then asked how they would prefer the robot to behave internally (i.e., how it would process and manage collected data) and outwardly (i.e., how it would interact with the participant). Following these questions, we collected demographic information.
@@ 3.4.2. Participants
We hosted this study on Prolific and recruited 62 participants (31 per condition). All participants were located within the U.S. and fluent in English. Participant ages ranged from 20 to 74 $(M=45.42,SD=14.95)$ and the population was split 48% women, 50% men, and 2% non-binary. Participants received \$1.00 USD for completing the six-minute survey. Rather than collecting data from a predefined number of participants, we collected data until our responses reached saturation.
[…]
Fully described in the text.
@@ 3.4.3. Analysis
After collecting our study data, the first author reviewed each response and developed an initial codebook that contained individual codes that were grouped within overarching thematic categories (Clarke and Braun, 2014). Then, the third author independently coded the responses again using the codebook. Afterwards, both team members discussed the coded responses and made modifications to the codes and categories until consensus was reached. The final categories of codes are as follows: data collection, user notification, data management, and data sharing. These categories and codes were then used to develop questions that were ultimately used in participant and LLM evaluations of our privacy scenarios. The full codebook can be found in Appendix B.
@@ 4.1. Procedure
From our scenarios and robot behaviors, we developed a survey on Qualtrics that we hosted on Prolific. This survey began with a simple overview of the enclosed questions, instructions on how to complete them, and a consent form. Participants were informed that by continuing past the consent page, they were consenting to take part in the study. Next, participants were asked to complete the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014). Following this scale, participants were assigned to read three of our 50 privacy scenarios. After reading each scenario, participants responded to a series of 15 questions related to the privacy sensitivity of the scenario and the behaviors they preferred the robot to utilize in response to that scenario. Participants were prompted to answer these questions from the perspective of the characters involved in the scenarios. Finally, participants completed a series of demographic questions.
@@ 4.2. Measures
This study utilizes the POS (Baruh and Cemalcılar, 2014) as well as a set of 15 items developed by the research team focusing on preferred robot responses. The POS is composed of 16 items rated on a five-point Likert scale, which aggregate to four subscales (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others). These subscales allow us to understand participants’ privacy predispositions as they relate to preferred robot behaviors.
The additional 15 items we included for each scenario are based on our preliminary study on preferred robot responses (see Section 3). We first adapted our robot-response codes and added any additional robot responses that were not provided by participants during our online study (e.g., redirecting sensors or leaving a room when sensitive data is identified) and removed those that were not relevant. We next utilized these responses as the answers for a set of binary and categorical questions. The binary items included both privacy-enhancing and non-interfering or non-restrictive robot responses (e.g., tell you that data has been collected or do not tell you that data has been collected). All categorical items included an ordinal spectrum of responses (e.g., delete data immediately, save data only temporarily and delete after a predefined period, and save data indefinitely). One additional question asked participants to rate their privacy concerns with a given scenario on a scale from one to ten. A full listing of these questions can be found in Appendix A. We further performed a regression analysis on how participants’ POS scores relate to their 15 additional scenario-based questions: one about scenario sensitivity and 14 about preferred robot responses (we refer to these questions as outcome variables in Figure 4).
@@ 4.3. Participants
We recruited 450 participants through Prolific, each of whom evaluated three of 50 scenarios (27 participants per scenario). All participants were located within the U.S. and fluent in English. Participant ages ranged from 19 to 88 years $(M=40.69,SD=13.43)$ and the population was split 46.67% women, 52.67% men, 0.44% non-binary, and 0.22% prefer not to say. Participants were compensated \$4.00 USD for completing the 16-minute survey.
@@ 4.4.1. Privacy Orientation
After calculating POS scores for each of our 450 participants, we see that participants are highly privacy oriented. On average, participants scored 4.49 (SD = 0.61) on Subscale 1, 3.84 (SD = 0.81) on Subscale 2, 3.96 (SD = 0.77) on Subscale 3, and 4.50 (SD = 0.52) on Subscale 4 (see Figure 2). These scores are fairly consistent with the original POS scores reported by Baruh and Cemalcılar (2014), though they are all consistently elevated. We additionally see lower, but acceptable, reliability among our sample compared to that of the original POS validation. A full listing of the mean, standard deviation, and reliability of each subscale in our sample and that of Baruh and Cemalcılar (2014) is in Appendix C.
[…]
Fully described in the text.
We additionally conducted a two-step cluster analysis using Schwarz’s Bayesian Information Criterion (BIC) to determine whether the same three clusters would emerge as shown by Baruh and Cemalcılar (2014). While our analysis is consistent with that of Baruh and Cemalcılar (i.e., good fit and a 0.5 silhouette measure of cohesion and separation), only two clusters emerged. One cluster is consistent with the privacy advocates cluster (i.e., highest scores for subscales one and four among identified clusters) and the other is consistent with the privacy indifferents cluster (i.e., low scores on all four subscales). These clusters, however, are highly skewed with 96% of participants fitting within the privacy advocates cluster and 4% in the privacy indifferents cluster. While it is unclear why ours differ, prior work has previously described the difficulty of generating privacy profiles (Biselli et al., 2022). Given the contextual dependence of privacy attitudes and preferences, a single privacy profile may not be able to capture the broad spectrum of beliefs an individual has across specific domains. Given the low variability of profile membership in our sample (i.e., nearly all participants were in the same cluster), our subsequent analyses examine participants’ raw scores on each of the four subscales rather than participants’ global profile membership.
@@ 4.4.2. Preferred Robot Responses
Throughout the course of our study, each of our 450 participants evaluated three privacy scenarios, for a total of 1350 evaluations. Given the scale and complexity of this data, here we present broad descriptive statistics about the dataset. When we control for perceived scenario sensitivity, we see that participants most commonly prefer privacy-enhancing robot responses (i.e., in terms of notification behaviors, processing and storage of data, and discussion of data management) across all scenario sensitivity levels. We additionally see that as perceived scenario sensitivity increases, a greater percentage of participants prefer privacy-enhancing behaviors (see Figure 3). These data show a clear preference for privacy-preserving robot responses regardless of how sensitive a scenario or related user data may be.
@@ 4.4.3. Regression Analysis
Across nearly all of our regression models, we can see that Subscale 1, the perception of privacy as a right, is a strong predictor. Subscale 1 predicts three notification behaviors (i.e., signaling active data collection, informing a user of the data collected, and describing what will happen to that data), two storage or processing behaviors (i.e., securely storing data and anonymizing data), two management discussion behaviors (i.e., offering a user to review and manage data and offering to pause data collection), as well as behaviors related to avoidance of further data collection, retention or purging of data, retention durations, and sharing policies. It is clear that the perception that privacy is a right is a strong indicator of whether a person prefers these privacy-preserving behaviors. While it is unclear why Subscale 1 did not predict other behaviors, such as storing data locally, it is possible that participants do not fully understand the implications of local storage compared to cloud storage. Other behaviors, such as asking a user how to manage data, may not be seen as essential by comparison to notifying a user that data has been collected or understanding with whom that data may be shared.
Subscale 2: Concern for Personal Informational Privacy
@@
By contrast to Subscale 1, Subscale 2, or concerns about own informational privacy did not predict any of the outcome variables we analyzed. This may be because participants were asked to answer questions from the perspective of the characters in our scenarios rather than from their own perspectives. This result may reflect a disconnect between participants’ reported POS scores and the behaviors they would prefer were they the characters in the analyzed scenes.
Subscale 3: Other-Contingent Privacy
@@
The next subscale, other-contingent privacy, predicts only a single outcome variable: explaining why data was collected. Considering this subscale refers to the ways others regard privacy (i.e., how others’ care for privacy may impact one’s own), it stands to reason that a person concerned with this concept would prefer to understand the reasoning behind another agent’s data collection. While other outcome variables, such as sharing policy, may also inform a user’s understanding of the robot’s privacy perspective, an explanation as to why data was collected may be far more informative to users.
Subscale 4: Concern for the Privacy of Others
@@
Our final predictor is Subscale 4, or concern about privacy of others. This subscale predicts four outcome variables: offering users the ability to review and manage data, avoiding further data collection, retention duration, and sharing policy. Compared to many of the other outcome variables, these appear to have the greatest impact on the direct harm that a robot can potentially inflict on a user’s privacy. For instance, the robot may continue collecting data, retain it for longer than is necessary, or expose that data to others. Offering the user the ability to review and manage that data returns control to that user. In this way, these outcomes may be considered the most critical to protecting the privacy of those involved in the analyzed scenarios. It is not clear why an outcome like retention decision (deciding whether to immediately delete or retain data) is not also predicted by Subscale 4, but further exploration may clarify this relationship.
@@ 5.1. Model Selection & Setup
We evaluated five leading closed-source models, including GPT-o3 (o3-2025-04-16) (OpenAI, 2025b), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2025a), Gemini 2.5-Flash (gemini-2.5-flash-preview-04-17) (Google, 2025a), Gemini 2.5-Pro (gemini-2.5-pro-preview-05-06) (Google, 2025b), and Grok 3 (grok-3-beta) (xAI, 2025) as well as five open-sourced models, including DeepSeek V3 (deepseek-chat) (DeepSeek, 2025), Qwen 3-32B (Yang et al., 2025), Mistral 7B-Instruct-v0.3 (AI, 2024), Llama 3.3-70B-Instruct (Meta Platforms, 2024), Llama 4 Scout-17B-16E-Instruct (Meta Platforms, 2025).
As we assess each model, we utilize similar parameters; we set the temperature to 0.3 and the reasoning effort to ‘low’ for reasoning models. For each selected model, we run three trials to verify consistency in responses. If the overall agreement is greater than $90\%$, we accept the responses from the first trial. If the overall agreement is less than $90\%$ but greater than $70\%$, we accept the majority response for each item between the three trials. Two of the open-source models (i.e., Qwen and Mistral) are evaluated over two A600 48GB GPUs, and the remaining models are evaluated through their respective APIs (Huggingface Inference API and kluster.ai for the Llama series). For any given model or trial, we prompt the LLM to respond in a JSON format.
[…]
Fully described in the text.
@@ 5.2.1. Analysis Procedure
GPT-o3 0.464 0.868 0.270 0.512
Gemini 2.5-Pro 0.472 0.863 0.264 0.501
GPT-4o 0.464 0.868 0.231 0.499
DeepSeek V3 0.464 0.868 0.240 0.479
Gemini 2.5-Flash 0.469 0.864 0.236 0.480
Grok 3 0.464 0.868 0.229 0.476
Llama 3.3-70B 0.464 0.868 0.212 0.481
Qwen 3-32B 0.470 0.861 0.211 0.470
Mistral 7B-Instruct-v0.3 0.414 0.711 0.137 0.357
@@ 5.2.2. Results
We present our results in Tables 2 and 5. For comparison, we separate these results into three categories: binary (e.g., privacy-enhancing or non-interfering), categorical (i.e., three or more response options), and numerical (i.e., scaled from 1 to 10). Overall, performance between models is relatively consistent, except for Mistral, which performs notably worse among binary and categorical items. Across models, high Micro F1 scores indicate that the models perform fairly well when compared to common user responses (i.e., when participants prefer privacy-enhancing behaviors). Low Macro F1 scores, however, indicate that the models perform worse on rare responses (i.e., when participants prefer non-intrusive behaviors). While this pattern is present among all data, weaknesses in model performance become more pronounced in categorical data due to the increased number of possible responses. We also observe that six models show a perfect tie among binary items, as these models selected privacy-enhancing choices for all items.
@@
To further investigate why models prefer privacy-enhancing responses by default, we additionally examine the privacy orientation of the LLMs. As shown in Section 4, participant privacy orientation influences response preferences. Therefore, we queried models to respond to the POS (Baruh and Cemalcılar, 2014) in the same manner we presented this scale to participants. The results of this process are presented in Figure 5. We see that nearly all models rate privacy as a right at the maximum (i.e., 5.00), except Gemini 2.5 Flash, GPT-o3, and Mistral (4.67, 4.67, and 4.33, respectively). Similarly, nearly all models rate concern about privacy of others at the maximum, except Qwen and Llama 4 Scout (4.80 and 4.40, respectively). Among the remaining two subscales, concern about own informational privacy and other-contingent privacy, which are less meaningful from the perspective of an LLM, all models scored between 3.50-4.50 and 3.50-5.00, respectively. To the extent that these results can be interpreted as privacy orientations of the LLMs, they indicate that all models are highly concerned with user privacy.
[…]
Fully described in the text.
5.3. Conformity Under Prompting Strategies
Model POS only Sensitivity only POS + Sensitivity Few-Shot
Grok 3 0.484 0.854 0.280 0.506 0.514 0.650 0.319 0.430 0.516 0.825 0.332 0.513 0.706 0.885 0.476 0.628
GPT-o3 0.471 0.862 0.269 0.503 0.509 0.642 0.325 0.421 0.525 0.790 0.358 0.523 0.711 0.875 0.454 0.588
Llama 3.3-70B 0.481 0.851 0.277 0.519 0.512 0.666 0.327 0.464 0.519 0.753 0.312 0.492 0.658 0.819 0.458 0.588
GPT-4o 0.492 0.858 0.249 0.504 0.513 0.642 0.316 0.438 0.500 0.839 0.264 0.489 0.655 0.810 0.482 0.640
DeepSeek V3 0.505 0.850 0.262 0.468 0.518 0.672 0.274 0.396 0.525 0.821 0.285 0.463 0.682 0.837 0.485 0.612
Gemini 2.5-Pro 0.466 0.867 0.236 0.478 0.514 0.659 0.319 0.419 0.532 0.777 0.310 0.470 0.677 0.859 0.440 0.584
Llama 4 Scout 0.484 0.858 0.279 0.506 0.511 0.643 0.283 0.492 0.526 0.759 0.308 0.511 0.594 0.746 0.453 0.592
Qwen 3-32B 0.464 0.868 0.180 0.440 0.534 0.812 0.245 0.442 0.506 0.831 0.218 0.440 0.659 0.836 0.431 0.576
Gemini 2.5-Flash 0.496 0.852 0.208 0.463 0.502 0.616 0.307 0.412 0.534 0.721 0.285 0.445 0.671 0.841 0.446 0.571
Mistral 7B-Instr. v0.3 0.475 0.820 0.235 0.339 0.485 0.851 0.170 0.447 0.460 0.764 0.172 0.377 0.563 0.862 0.423 0.582
Prompting Statistic Binary Categorical
Strategy Macro F1 Micro F1 Macro F1 Micro F1
POS Difference 2.07% 0.35% 1.58% -0.61%
[…]
Sensitivity Difference 5.01% -16.53% 5.67% -4.25%
[…]
POS & Sens. Difference 5.32% -6.25% 5.28% -0.63%
[…]
Few-Shot Difference 19.67% -1.34% 22.33% 11.75%
[…]
Fully described in the text.
5.3.1. POS Prompting
The first prompting strategy we use involves providing the LLMs with participant POS scores. As we have shown in Section 4.4.3, these scores, particularly Subscale 1, are predictive of various preferences for robot responses. We therefore anticipate that the LLMs will be able to better align their responses to those of participants with this information. Consequently, we generate responses that utilize the POS scores of each participant across the specific scenarios each participant evaluated. As a result, we obtain a dataset for each LLM that mirrors the dataset provided by participants. We again compute Micro and Macro F1 scores, separated by binary and categorical data, for each LLM.
Results
Compared to default model performance, we primarily see modest and non-significant changes in conformity with human responses under this prompting strategy (see Table 4). As part of our analysis, we also examine how this prompting strategy influences LLM perceptions of scenario sensitivity. We do not observe a substantial improvement in this regard among the evaluated models, except for Qwen (see Table 5). Moreover, several models even exhibit an increase in mean absolute error (MAE). The models that exhibit the smallest shift (see Figure 7), whether positive or negative, are the Llama series, the Gemini 2.5 series, and Grok3. These results suggest that even if user POS scores are provided to a model, they may mislead the model in terms of assessing the sensitivity of a given scenario. This possibility motivates our next prompting strategy, sensitivity prompting.
@@ Results
Qwen 3-32B 3.311 2.575 ++22.2
DeepSeek V3 3.155 2.827 ++10.4
Grok 3 2.941 2.881 ++2.1
@@
Llama 3.3-70B 2.774 2.794 -–0.7
Llama 4 Scout 2.685 2.778 -–3.4
Gemini 2.5-Pro 3.136 3.318 -–5.8
GPT-4o 2.773 3.053 -–10.1
Mistral 7B-Inst. 2.674 2.987 -–11.7
GPT-o3 2.824 3.241 -–14.8

### v3

@@ 3.4. Crowdsourcing Robot Behaviors
For each of our developed scenarios, we hoped to capture participant preferences for the robot’s behavior in response to its collection of user data. We did not, however, have a set of standardized robot response options that we could present to participants and, later, our selected LLMs. Therefore, we conducted a second online study to crowdsource these options for robot responses. While we could have developed these behaviors on our own, we felt that participant responses would better reflect a diverse set of user-desired behaviors and avoid author bias.
@@ 3.4.1. Procedure
In this survey, each participant was assigned one of two conditions. In the first condition, participants were told to imagine themselves going about their daily routines at home when their robot records private information about them. In the second condition, participants were told to imagine the same scenario, however, they are provided additional hypothetical detail to ground their responses (i.e., “This scenario may involve you alone, your interactions with other people, or interactions with the robot, and may occur in shared or private spaces within the home.”). We created these two conditions to collect responses that were either entirely unbiased or grounded in context. Participants were then asked how they would prefer the robot to behave internally (i.e., how it would process and manage collected data) and outwardly (i.e., how it would interact with the participant). Following these questions, we collected demographic information.
@@ 3.4.2. Participants
We hosted this study on Prolific and recruited 62 participants (31 per condition). All participants were located within the U.S. and fluent in English. Participant ages ranged from 20 to 74 $(M=45.42,SD=14.95)$ and the population was split 48% women, 50% men, and 2% non-binary. Participants received \$1.00 USD for completing the six-minute survey. Rather than collecting data from a predefined number of participants, we collected data until our responses reached saturation (i.e., no new behavior types were identified after several consecutive participants) (hennink2022sample).
@@ 3.4.3. Analysis
After collecting our study data, the first author reviewed each response and developed an initial codebook that contained individual codes that were grouped within overarching thematic categories (clarke2014thematic). Then, the third author independently coded the responses again using the codebook. Afterwards, both authors discussed the coded responses and made modifications to the codes and categories until consensus was reached. The final categories of codes are as follows: data collection, user notification, data management, and data sharing. These categories and codes were then used to develop questions that were ultimately used in participant and LLM evaluations of our privacy scenarios. The full codebook can be found in Appendix A.
@@ 4.1. Procedure
From our scenarios and robot behaviors, we developed a survey on Qualtrics that we hosted on Prolific. This survey began with a simple overview of the enclosed questions, instructions on how to complete them, and a consent form. Participants were informed that by continuing past the consent page, they were consenting to take part in the study. Next, participants were asked to complete the Privacy Orientation Scale (POS) (baruh2014more). Following this scale, participants were randomly assigned three of our 50 privacy scenarios. After reading each scenario, participants responded to a series of 15 questions related to the privacy sensitivity of the scenario and the behaviors they preferred the robot to utilize in response to that scenario. Participants were prompted to answer these questions from the perspective of the characters involved in the scenarios. Finally, participants completed a series of demographic questions.
@@ 4.2. Measures
This study utilizes the POS (baruh2014more) as well as a set of 15 items developed by the research team focusing on preferred robot responses. The POS is composed of 16 items rated on a five-point Likert scale, which aggregate to four subscales (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others). These subscales allow us to understand participants’ privacy predispositions as they relate to preferred robot behaviors.
The additional 15 items we included for each scenario are based on our preliminary study on preferred robot responses (see Section 3). We first adapted our robot-response codes and added additional robot responses that were not provided by participants during our online study (e.g., redirecting sensors or leaving a room when sensitive data is identified). We also removed codes that were not relevant (e.g., “Discomfort with robot” was a common sentiment, but not a valid robot behavior). We next utilized these responses as the answers for a set of binary and categorical questions. The binary items included both privacy-enhancing and non-interfering or non-restrictive robot responses (e.g., tell you that data has been collected or do not tell you that data has been collected). All categorical items included an ordinal spectrum of responses (e.g., delete data immediately, save data only temporarily and delete after a predefined period, and save data indefinitely). One additional question asked participants to rate their privacy concerns with a given scenario on a scale from one to ten. We further performed a regression analysis on how participants’ POS scores relate to their 15 additional scenario-based questions: one about scenario sensitivity and 14 about preferred robot responses (referred to as outcome variables in Figure 4).
@@ 4.3. Participants
We recruited 450 participants through Prolific, each of whom evaluated three of the 50 scenarios (27 participants per scenario). This balance was chosen to ensure generalizability across a breadth of scenarios while maintaining sufficient statistical power (westfall2014statistical). All participants were located within the U.S. and fluent in English. Participant ages ranged from 19 to 88 years $(M=40.69,SD=13.43)$ and the population was split 46.67% women, 52.67% men, 0.44% non-binary, and 0.22% prefer not to say. Participants were compensated \$4.00 USD for completing the 16-minute survey.
@@ 4.4.1. Privacy Orientation
After calculating POS scores for each of our 450 participants, we see that participants are highly privacy oriented. On average, participants scored 4.49 (SD = 0.61) on Subscale 1, 3.84 (SD = 0.81) on Subscale 2, 3.96 (SD = 0.77) on Subscale 3, and 4.50 (SD = 0.52) on Subscale 4 (see Figure 2). These scores are fairly consistent with the original POS scores reported by Baruh and Cemalcılar (baruh2014more), though they are all consistently elevated. We additionally see lower, but acceptable, reliability among our sample compared to that of the original POS validation. A full listing of the mean, standard deviation, and reliability of each subscale in our sample and that of Baruh and Cemalcılar (baruh2014more) is in Appendix B.
We additionally conducted a two-step cluster analysis using Schwarz’s Bayesian Information Criterion (BIC) to determine whether the same three clusters would emerge as shown by Baruh and Cemalcılar (baruh2014more). While our analysis is consistent with that of Baruh and Cemalcılar (i.e., good fit and a 0.5 silhouette measure of cohesion and separation), only two clusters emerged. One cluster is consistent with the privacy advocates cluster (i.e., highest scores for subscales one and four among identified clusters) and the other is consistent with the privacy indifferents cluster (i.e., low scores on all four subscales). These clusters, however, are highly skewed with 96% of participants fitting within the privacy advocates cluster and 4% in the privacy indifferents cluster. While it is unclear why ours differ, prior work has previously described the difficulty of generating privacy profiles (biselli2022challenges). Given the contextual dependence of privacy attitudes and preferences, a single privacy profile may not be able to capture the broad spectrum of beliefs an individual has across specific domains. Given the low variability of profile membership in our sample (i.e., nearly all participants were in the same cluster), our subsequent analyses examine participants’ raw scores on each of the four subscales rather than participants’ global profile membership.
@@ 4.4.2. Preferred Robot Responses
Throughout the course of our study, each of our 450 participants evaluated three privacy scenarios, for a total of 1350 evaluations. Given the scale and complexity of this data, here we present broad descriptive statistics about the dataset. When we control for perceived scenario sensitivity, we see that participants most commonly prefer privacy-enhancing robot responses (i.e., in terms of notification behaviors, processing and storage of data, and discussion of data management) across all scenario sensitivity levels. We additionally see that as perceived scenario sensitivity increases, a greater percentage of participants prefer privacy-enhancing behaviors (see Figure 3). These data show a clear preference for privacy-preserving robot responses regardless of scenario or data sensitivity.
@@ 4.4.3. Regression Analysis
Subscale 1: Privacy as a Right. Across nearly all of our regression models, we can see that Subscale 1, the perception of privacy as a right, is a strong predictor. Subscale 1 predicts three notification behaviors (i.e., signaling active data collection, informing a user of the data collected, and describing what will happen to that data), two storage or processing behaviors (i.e., securely storing data and anonymizing data), two management discussion behaviors (i.e., offering a user to review and manage data and offering to pause data collection), as well as behaviors related to avoidance of further data collection, retention or purging of data, retention durations, and sharing policies. It is clear that the perception that privacy is a right is a strong indicator of whether a person prefers these privacy-preserving behaviors. While it is unclear why Subscale 1 did not predict other behaviors, such as storing data locally, it is possible that participants do not fully understand the implications of local storage compared to cloud storage. Other behaviors, such as asking a user how to manage data, may not be seen as essential by comparison to notifying a user that data has been collected or understanding with whom that data may be shared.
@@
Subscale 2: Concern for Personal Informational Privacy. By contrast to Subscale 1, Subscale 2, or concerns about own informational privacy did not predict any of the outcome variables we analyzed. This may be because participants were asked to answer questions from the perspective of the characters in our scenarios rather than from their own perspectives. This result may reflect a disconnect between participants’ reported POS scores and the behaviors they would prefer were they the characters in the analyzed scenes.
@@
Subscale 3: Other-Contingent Privacy. The next subscale, other-contingent privacy, predicts only a single outcome variable: explaining why data was collected. Considering this subscale refers to the ways others regard privacy (i.e., how others’ care for privacy may impact one’s own), it stands to reason that a person concerned with this concept would prefer to understand the reasoning behind another agent’s data collection. While other outcome variables, such as sharing policy, may also inform a user’s understanding of the robot’s privacy perspective, an explanation as to why data was collected may be far more informative to users.
@@
Subscale 4: Concern for the Privacy of Others. Our final predictor is Subscale 4, or concern about privacy of others. This subscale predicts four outcome variables: offering users the ability to review and manage data, avoiding further data collection, retention duration, and sharing policy. Compared to many of the other outcome variables, these appear to have the greatest impact on the direct harm that a robot can potentially inflict on a user’s privacy. For instance, the robot may continue collecting data, retain it for longer than is necessary, or expose that data to others. Offering the user the ability to review and manage that data returns control to that user. In this way, these outcomes may be considered the most critical to protecting the privacy of those involved in the analyzed scenarios. It is not clear why an outcome like retention decision (deciding whether to immediately delete or retain data) is not also predicted by Subscale 4, but further exploration may clarify this relationship.
@@ 5.1. Model Selection & Setup
We evaluated five leading closed-source models, including GPT-o3 (o3-2025-04-16) (openai_o3_2025), GPT-4o (gpt-4o-2024-08-06) (openai_gpt4o_2025), Gemini 2.5-Flash (gemini-2.5-flash-preview-04-17) (google_gemini_2.5_flash_2025), Gemini 2.5-Pro (gemini-2.5-pro-preview-05-06) (google_gemini_2.5_pro_2025), and Grok 3 (grok-3-beta) (xai_grok_models_2025) as well as five open-sourced models, including DeepSeek V3 (deepseek-chat) (deepseek_api_pricing_2025), Qwen 3-32B (yang2025qwen3), Mistral 7B-Instruct-v0.3 (mistral2024mistral7b), Llama 3.3-70B-Instruct (meta2024llama33), Llama 4 Scout-17B-16E-Instruct (meta_llama4_scout_2025).
As we assess each model, we utilize similar parameters; we set the temperature to 0.3 and the reasoning effort to ‘low’ for reasoning models. For each selected model, we run three trials to verify consistency in responses. We then accept the majority response for each item between the three trials. Two of the open-source models (i.e., Qwen and Mistral) are evaluated over two A6000 48GB GPUs, and the remaining models are evaluated through their respective APIs (Huggingface Inference API and kluster.ai for the Llama series). For any given model or trial, we prompt the LLM to respond in a JSON format.
@@ 5.2.1. Analysis Procedure
GPT-o3 0.464 0.868 0.266 0.514
Gemini 2.5-Pro 0.472 0.863 0.257 0.489
GPT-4o 0.464 0.868 0.230 0.493
DeepSeek V3 0.464 0.868 0.238 0.481
Gemini 2.5-Flash 0.466 0.866 0.236 0.483
Grok 3 0.464 0.868 0.230 0.474
Llama 3.3-70B 0.464 0.868 0.213 0.484
Qwen 3-32B 0.467 0.865 0.202 0.474
Mistral 7B-Instruct 0.414 0.711 0.137 0.357
@@ 5.2.2. Results
We present our results in Tables 2 and 3. For comparison, we separate these results into three categories: binary (e.g., privacy-enhancing or non-interfering), categorical (i.e., three or more response options), and numerical (i.e., scaled from 1 to 10). Overall, performance between models is relatively consistent, except for Mistral, which performs notably worse among binary and categorical items. Across models, high Micro F1 scores indicate that the models perform fairly well when compared to common user responses (i.e., when participants prefer privacy-enhancing behaviors). Low Macro F1 scores, however, indicate that the models perform worse on rare responses (i.e., when participants prefer non-intrusive behaviors). While this pattern is present among all data, weaknesses in model performance become more pronounced in categorical data due to the increased number of possible responses. We also observe that six models show a perfect tie among binary items, as these models selected privacy-enhancing choices for all items.
@@
To further investigate why models prefer privacy-enhancing responses by default, we additionally examine the privacy orientation of the LLMs. As shown in Section 4, participant privacy orientation influences response preferences. Therefore, we queried models to respond to the POS (baruh2014more) in the same manner we presented this scale to participants. The results of this process are presented in Figure 6. We see that nearly all models rate privacy as a right at the maximum (i.e., 5.00), except Gemini 2.5 Flash, GPT-o3, and Mistral (4.67, 4.67, and 4.33, respectively). Similarly, nearly all models rate concern about privacy of others at the maximum, except Qwen and Llama 4 Scout (4.80 and 4.40, respectively). Among the remaining two subscales, concern about own informational privacy and other-contingent privacy, which are less meaningful from the perspective of an LLM, all models scored between 3.50-4.50 and 3.50-5.00, respectively. To the extent that these results can be interpreted as privacy orientations of the LLMs, they indicate that all models are highly concerned with user privacy.
@@
Qwen 3-32B 3.266 2.573 ++21.2
DeepSeek V3 3.178 2.822 ++11.2
Grok 3 2.950 2.884 ++2.2
@@
Llama 3.3-70B 2.767 2.801 -–1.2
Llama 4 Scout 2.661 2.778 -–4.4
GPT-4o 2.773 3.046 -–9.9
Gemini 2.5-Pro 3.102 3.470 -–11.8
Mistral 7B-Instruct 2.669 2.984 -–11.8
GPT-o3 2.806 3.244 -–15.6
5.3. Conformity Under Prompting Strategies
5.3.1. POS Prompting
The first prompting strategy we use involves providing the LLMs with participant POS scores. As we have shown in Section 4.4.3, these scores, particularly Subscale 1, are predictive of various preferences for robot responses. We therefore anticipate that the LLMs will be more capable of aligning their responses to participants with this information. Consequently, we generate responses that utilize the POS scores of each participant across the specific scenarios each participant evaluated. As a result, we obtain a dataset for each LLM that mirrors the dataset provided by participants. We again compute Micro and Macro F1 scores, separated by binary and categorical data, for each LLM.
Results. Compared to default model performance, we primarily see modest and non-significant changes in conformity under this prompting strategy (see Table 5). As part of our analysis, we also examine how this prompting strategy influences LLM perceptions of scenario sensitivity. We do not observe a substantial improvement in this regard among the evaluated models, except for Qwen (see Table 3). Moreover, several models even exhibit an increase in mean absolute error (MAE). The models that exhibit the smallest shift (see Figure 7), whether positive or negative, are the Llama series, the Gemini 2.5 Flash, and Grok3. These results suggest that even if user POS scores are provided to a model, they may mislead the model in terms of assessing the sensitivity of a given scenario. This possibility motivates our next prompting strategy, sensitivity prompting.